Answer with all steps and explaintion please:

O Are these series convergent or divengont? which test did you use? (A À { (-1) sin(n) B
Answer #1

Infinite Calculator: Given, (-1) sin(n) P=1 We have to find the convergence or divergence of the given series. Since, lim (-Since, - In x Inx =lim 100 ch . 1 d =lim (In x) -0.2 dx =lim (in x) [com =lim (int 100 = lim (Int) 100 = 0 Therefore, by the